# Optimal Parallel $\lambda$-Reduction

### v2

@@ Abstract
I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the model and the method constitute an algorithm for optimal parallel $\lambda$-reduction, solving the longstanding enigma with groundbreaking clarity. I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets—one that severely restricts the structure of sharing, among other drawbacks. Unhindered by these restrictions, the $\Delta$-Nets model opens the door to new highly parallel programming language implementations and computer architectures that are more efficient and performant than previously possible.
@@ 1.  Introduction
For example, when reducing the non-normalizing $\lambda$-term $(\lambda x.x\,x)(\lambda y.y\,y)$ with these algorithms, every iteration creates new delimiters. Each new delimiter takes up additional memory (space), and leads to additional unnecessary interactions (time). Additionally, the existing algorithms fail to establish a global partial ordering, which is absolutely critical to ensure that all nets associated with normalizing $\lambda$-terms normalize. Not only are the existing algorithms unsatisfactory from a theoretical perspective, their inefficiencies preclude them from being used at the core of programming language implementations.
In this paper, I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the system and translation method constitute an algorithm for optimal parallel $\lambda$-reduction. As an interaction system, the core of the model is perfectly confluent: every normalizing reduction order produces the same result in the same number of steps. The core model is then necessarily extended with non-interaction canonicalization rules, which depart from the interaction paradigm. In order to minimize total number of reductions and guarantee normalization for all nets associated with normalizing $\lambda$-terms, a global partial reduction ordering needs to be followed, and is specified. The $\Delta$-Nets algorithm solves the longstanding enigma of optimal $\lambda$-calculi reduction with groundbreaking clarity. Instead of making use of delimiters, sharing is expressed through a single agent type which allows any number of auxiliary ports, called a replicator. Each instance of a replicator incorporates information that in previous models was spread across multiple agents, such as indexed fans and delimiters. This consolidation of information enables simplifications that were previously unfeasible, and leads to constant memory usage in the reduction of $(\lambda x.x\,x)(\lambda y.y\,y)$, for example. Finally, I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets. The additional degrees of freedom in $\Delta$-Nets allow it to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is performed which is rendered unneccessary later, and no reduction operation which is necessary is performed more than once.
@@ 3.  From $\lambda$-terms to $\Delta$-nets
Note that both applications and abstractions are represented by fans, and $\beta$-reduction is expressed through fan annihilation. Whether a specific fan in a $\Delta$-net represents an abstraction or an application can always be determined by inspecting the net (without fan labels) and tracing paths from the root node. While abstraction fans have two parent ports and one child port, application fans have one parent port and two child ports. Every port of any node (including non-agent nodes) in a proper $\Delta$-net is either a parent port or a child port, and every wire connects a parent port with a child port. Although this duality has no direct bearing on the interaction process, it is clearly present, and it is instrumental in some discussions.
@@ 4.  From $\Delta$-nets back to $\lambda$-terms
The only interaction rule in $\Delta L$-Nets is fan annihilation, which expresses $\beta$-reduction. Therefore if $n$ $\beta$-reductions normalize a $\lambda L$-term $t$, then $n$ interactions normalize the $\lambda L$-net $\phi_{L}(t)$. As a result, in the $\Delta L$-Nets system, all proper nets are canonical, and interactions can be applied in any order, with perfect confluence. Compared to $\Delta L$-Nets, the other $\Delta$-Nets systems have additional interaction rules, which don’t have $\lambda$-calculus analogues.
The $\Delta A$-Nets system includes three core interaction rules: fan annihilation, fan erasure, and eraser annihilation. While in the $\lambda$-calculus a single $\beta$-reduction can erase a term of arbitrary size, erasure in $\Delta$-Nets is a computational process whose space and time complexity depends, respectively, on the size and structure of the net being erased. This makes $\Delta$-reductions a more useful measure of computational complexity than $\beta$-reductions.
Applying an abstraction which doesn’t use its bound variable results in an eraser becoming connected to a parent port. Such an eraser can erase abstraction fans and fan-out replicators, but when it reaches the parent port of an application fan, for example, which is an auxiliary port, the erasure process stops. As a result, absent additional rules, irreducible subnets may be produced which don’t have a $\lambda$-calculus analogue. For example, applying an abstraction which doesn’t use its bound variable to an argument which only uses globally-free variables produces a subnet which is disjointed from the root. In order to eliminate all such subnets, two canonicalization rules are introduced: fan decay and free-variable erasure. These rules are illustrated in Figure 6.
$fv$
Figure 6: The $\Delta A$-Nets canonicalization rules:
fan decay and free-variable erasure.
Fan decay allows the erasure of application fans via their parent ports and also of abstraction fans via their body ports. Free-variable erasure allows the erasure of unused free variables. The fan decay rule does not need to differentiate between abstraction and application fans. If an eraser becomes connected to the body port of an abstraction fan, there will be a counterpart eraser somewhere above in the net, connected to a parent port, such that it is safe to decay the abstraction fan.
In the $\Delta A$-Nets system, all three core interaction rules and the two canonicalization rules can be applied in any order, always producing the same result. However, in order to minimize the total number of reduction steps, the fan decay rule should only be applied when the fan annihilation rule cannot be applied. Since in this system only the application of the fan annihilation rule can lead to the creation of new fan-fan active pairs, this partial ordering ensures that all fan annihilations are performed before all fan decays. When no further reduction operation can be applied and the net is normal, it becomes a canonical $\Delta A$-net.
@@
During reduction of $\Delta I$- and $\Delta K$-Nets, replicators can combine into replicator trees. A replicator tree is a subnet containing only replicator agents such that each one’s principal port is connected to an auxiliary port of another, except for the tree’s root replicator. In general, replicators that belong to the same tree cannot be merged during reduction because they may be paired. However, if consecutive replicators in a replicator tree are known to be unpaired, they can be safely merged together. Unpaired replicator merging is a canonicalization rule present in both $\Delta I$- and $\Delta K$-Nets. Additionally, in the $\Delta K$-Nets system, it is possible to eliminate unpaired replicators’ auxiliary ports which are connected to erasers (partial unpaired replicator decay). If all auxiliary ports of a replicator are connected to erasers, the entire replicator (regardless of its status) is erased and replaced by an eraser (complete replicator decay). If, as part of partial unpaired replicator decay, an unpaired replicator is left with a single auxiliary port with a level delta of zero, it is replaced by a wire, as it is equivalent to one. These replicator decay rules are $\Delta K$-Nets canonicalization rules.
@@
Replicators start as unpaired, and this status can be propagated across interactions. When an unpaired replicator interacts with a fan, the status changes and the two resulting replicators are no longer unpaired. If an unpaired replicator (A) is connected to a consecutive replicator (B) of unknown status via an auxiliary port, and a certain local constraint is met, then the consecutive replicator can be determined to be unpaired, and the two can then be merged. The constraint is met when the second replicator’s level is greater than or equal to the first replicator’s level, but no greater than the first replicator’s level plus the level delta of the auxiliary port that connects them: $0\leq l_{B}-l_{A}\leq d$. Under this constraint, no replicator is able to interact with the second replicator before the first replicator is annihilated. Since the first replicator is unpaired, it can never be annihilated, and the second one must be unpaired as well.
Before applying a rule that involves an unpaired replicator (other than replicator erasure), it is always best to merge the unpaired replicator with consecutive unpaired replicators and decay it if possible. Relative to a baseline reduction order in which replicators are only merged when no other reductions are possible, this can only reduce the total number of reductions and the total number of agents, improving space and time efficiency. Unpaired replicators can be merged lazily, immediately before the fan replication, replicator replication, and aux fan replication rules. When multiple rules can be applied, it is advantageous to apply rules involving unpaired replicators last (specially those with two unpaired replicators) as this allows replicators of unknown status to accumulate behind unpaired replicators and potentially be determined to be unpaired and then merged.
@@
It is beneficial to limit replicators to have at most two auxiliary ports. This imposes the smallest possible upper bound on interaction rule complexity and agent size, making the total number of interactions an effective measure of time complexity, and the total number of agents an effective measure of space complexity. A tree of such replicators can stand in for any replicator with any number of auxiliary ports. There are many ways of building such trees, each with different trade-offs.
@@
The $\Delta I$-Nets system includes all core interaction rules which involve fans and replicators and don’t involve erasers. Additionally, it includes two canonicalization rules: unpaired replicator merging and aux fan replication (illustrated in Figure 7). The $\Delta I$-Nets core interaction rules with lazy unpaired replicator merging, and the aux fan replication canonicalization rule can be applied in any order, always producing the same result. However, in order to minimize the total number of reduction steps, the reduction process needs to be split in two phases. In the first phase the $\Delta I$-Nets core rules with lazy unpaired replicator merging are applied in any order until no active pair is left. The reduction process then switches to the second phase, in which the aux fan replication rule replaces the core fan replication rule. This second phase can be alternatively understood as modifying all fans such that their first auxiliary port becomes the principal port. Similarly to the fan decay rule, the aux fan replication rule does not need to differentiate between abstraction and application fans. If a replicator becomes connected to the body port of an abstraction fan, there will be counterpart replicators positioned such that it is safe to aux-replicate the abstraction fan. This process transforms the sharing structures such that all fan-out replicators are eliminated, all appropriate subnets are replicated, and all fan-in replicators accumulate at the variable port of abstraction fans. Finally, consecutive replicators can be merged since they are all unpaired, resulting in a canonical $\Delta I$-net.
The $\Delta K$-Nets system includes all core interaction rules (augmented with lazy unpaired replicator merging), and all the canonicalization rules that are part of $\Delta A$-Nets and $\Delta I$-Nets. Additionally it includes the replicator decay rules mentioned previously. A complex partial ordering of reduction operations is required in order to minimize the total number of reduction steps and, more importantly, to ensure that all $\Delta K$-nets associated with normalizing $\lambda K$-terms normalize. First, all core annihilations and erasures, including free variable erasure, are applied in any order. When no fan annihilations are possible, fan decays can be applied. Separately, when no replicator annihilations are possible, replicator decays can be applied. This ensures that fans and replicators are only eliminated through decay rules when they cannot be annihilated, which is more efficient. When no such reductions are possible (with the exception of eraser annihilation and free variable erasure, which can be applied at any time), core commutations are performed, augmented with lazy unpaired replicator merging. Commutations can only be applied when no annihilations, erasures, and decays are possible, otherwise, replications could be performed in subnets undergoing erasure, resulting, in the best case, in suboptimal behavior, and in the worst case in non-termination when termination is possible. Among the commutations, active pairs with unpaired replicators should be reduced last, as previously mentioned, in order to maximize the number of replicators that can be merged. When no commutations are possible, the aux fan replication rule is applied. Note that these are not separate phases, as in the $\Delta I$-Nets system, but rather a global partial ordering. Finally, when no further reduction operations can be applied, consecutive replicators can be merged since they are all unpaired, resulting in a canonical $\Delta K$-net.
Since all normal $\Delta$-nets are canonical, the $\Delta$-Nets systems are all confluent and possess the Church-Rosser property. The partial ordering of reduction operations specified for $\Delta K$-Nets is optimal in all $\Delta S$-Nets systems, $\forall S\in\Sigma$, omitting the appropriate reduction operations in each case. For ease of reference, the optimal $\Delta$-Nets partial order is summarized below. With the exception of reduction operations in group 0, which can be applied at any time, reduction operations in each numbered group can only be applied if no reduction operations in lower-numbered groups are possible. An asterisk indicates that a reduction operation is not a core interaction rule but a canonicalization rule instead. The systems to which each reduction operation applies are indicated in square brackets.
  1. 0.
    Unordered
    • $\bullet$
      Eraser annihilation [A,K]
    • $\bullet$
      Free-variable erasure [A,K]*
  2. 1.
    Deletions
    1. (a)
      Annihilations
      • $\bullet$
        Fan annihilation [L,A,I,K]
      • $\bullet$
        Replicator annihilation [I,K]
    2. (b)
      Erasures
      • $\bullet$
        Fan erasure [A,K]
      • $\bullet$
        Replicator erasure [K]
    3. (c)
      Decays
      Each should only be applied when no annihilations of the same agent type are possible, as annihilations delete agents in less interactions.
      • $\bullet$
        Fan decay [A,K]*
      • $\bullet$
        Complete replicator decay [K]*
  3. 2.
    Commutations
    • $\bullet$
      Fan replication [I,K]
    • $\bullet$
      Replicator replication [I,K]
  4. 3.
    Expansions
    • $\bullet$
      Aux fan replication [I,K]*
  5. 4.
    Replicator Canonicalizations
    At this stage, only unpaired fan-in replicators remain, therefore all replicator trees are merged and (partially) decayed.
    • $\bullet$
      Unpaired replicator merging [I,K]*
    • $\bullet$
      Partial unpaired replicator decay [K]*
Before applying fan replication, replicator replication, or aux fan replication, each involved unpaired replicator should:
  • $\bullet$
    be merged with all unpaired replicators which are connected to its auxiliary ports (unpaired replicator merging) [I,K]*
  • $\bullet$
    have all auxiliary ports connected to erasers removed (partial unpaired replicator decay) [K]*
@@ 5.  Conclusion
The existence of non-linear non-canonical proper $\Delta$-nets reflects the additional degrees of freedom that gradual erasure and interior sharing introduce, which are not present in the $\lambda$-calculus. A given $\lambda A$-term, for example, can potentially be represented by many different proper $\Delta A$-nets, which differ only with respect to the progress of gradual erasure. Similarly, a given $\lambda I$-term can potentially be represented by many different proper $\Delta I$-nets, which differ only with respect to their sharing structure. As an example, take a proper $\lambda I$-net $n$, with $\phi_{I}^{-1}(\Omega_{I}(n))=MM$, for some $\lambda I$-term $M$. The $\lambda I$-net $n$ may have two distinct but equal subnets that each represent $M$, or it may have a single such subnet which is shared among the two occurrances in the application. In fact, the sharing structure of $n$ could be arbitrarily complex. These additional degrees of freedom allow $\Delta$-Nets to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is performed which is rendered unneccessary later, and no reduction operation which is necessary is performed more than once.

### v3

@@ Abstract
I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the model and the method constitute an algorithm for optimal parallel $\lambda$-reduction, solving the longstanding enigma with groundbreaking clarity. I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets—one that severely restricts the structure of sharing, among other drawbacks. Unhindered by these restrictions, the $\Delta$-Nets model opens the door to new parallel programming language implementations and computer architectures that are more efficient and performant than previously possible.
@@ 1.  Introduction
For example, when reducing the non-normalizing $\lambda$-term $(\lambda x.x\,x)(\lambda y.y\,y)$ with these algorithms, every iteration creates new delimiters. Each new delimiter takes up additional memory (space), and leads to additional unnecessary interactions (time). Additionally, the existing algorithms fail to establish a global reduction order, which is unfortunately critical to ensure that all nets associated with normalizing $\lambda$-terms normalize. Not only are the existing algorithms unsatisfactory from a theoretical perspective, their inefficiencies preclude them from being used at the core of programming language implementations.
In this paper, I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the system and translation method constitute an algorithm for optimal parallel $\lambda$-reduction. As an interaction system, the core of the model is perfectly confluent: every normalizing reduction order produces the same result in the same number of steps. The core model is then necessarily extended with non-interaction canonicalization rules, which depart from the interaction paradigm. Along with a global reduction order, canonicalizations ensure Church–Rosser confluence and optimality in nonlinear systems. The $\Delta$-Nets algorithm solves the longstanding enigma of optimal $\lambda$-calculi reduction with groundbreaking clarity. Instead of making use of delimiters, sharing is expressed through a single agent type which allows any number of auxiliary ports, called a replicator. Each instance of a replicator incorporates information that in previous models was spread across multiple agents, such as indexed fans and delimiters. This consolidation of information enables simplifications that were previously unfeasible, and leads to constant memory usage in the reduction of $(\lambda x.x\,x)(\lambda y.y\,y)$, for example. Finally, I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets. The additional degrees of freedom in $\Delta$-Nets allow it to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is applied which is rendered unneccessary later, and no reduction operation which is necessary is applied more than once.
@@ 3.  From $\lambda$-terms to $\Delta$-nets
Note that both applications and abstractions are represented by fans, and $\beta$-reduction is expressed through fan annihilation. Whether a specific fan in a $\Delta$-net represents an abstraction or an application can always be determined by inspecting the net (without fan labels) and tracing paths from the root node. While abstraction fans have two parent ports and one child port, application fans have one parent port and two child ports. Every port of any node (including non-agent nodes) in a proper $\Delta$-net is either a child port or a parent port, and every wire connects a child port with a parent port. Although this duality has no direct bearing on the interaction process, it is clearly present, and it is instrumental in some discussions.
@@
In some ways, a fan is just a replicator with two auxiliary ports, zero level, and zero deltas. If replicators were allowed to have zero auxiliary ports, then an eraser could also just be a replicator with no auxiliary ports and zero level (or no level). It’s tempting to attempt to consolidate the three agents into one—the $\Delta$-agent, if you will. However, fans and replicators have a critical distinction in the types of their auxiliary ports: a replicator’s auxiliary ports are either all parent ports or all child ports, while a fan has one of each. This affects how replicator pairedness is tracked: after fan replication the resulting replicators become paired whereas in replicator replication they keep their original status.
In standard $\Delta$-Nets, replicators have absolute levels, i.e., the initial level of each replicator is exactly determined by how many times a path from the root to the replicator’s principal port traverses fans out of their second auxiliary ports. There exists a dual formulation of the $\Delta$-Nets system to this where replicators have relative levels—all replicators start with level zero, and a replicator’s level gets incremented when it traverses out of the second auxiliary port of a fan.
@@ 4.  From $\Delta$-nets back to $\lambda$-terms
The only interaction rule in $\Delta L$-Nets is fan annihilation, which expresses $\beta$-reduction. Therefore if $n$ $\beta$-reductions normalize a $\lambda L$-term $t$, then $n$ interactions normalize the $\lambda L$-net $\phi_{L}(t)$. As a result, in the $\Delta L$-Nets system, all proper nets are canonical, and fan annihilations can be applied in any order, with perfect confluence. Compared to $\Delta L$-Nets, the other $\Delta$-Nets systems have additional interaction rules, which don’t have $\lambda$-calculus analogues.
In $\Delta$-Nets systems with erasure, applying an abstraction which doesn’t use its bound variable results in an eraser becoming connected to a parent port. Such an eraser could erase abstraction fans and fan-out replicators, but if it reaches the parent port of an application fan, for example, which is an auxiliary port, the erasure process would cease. As a result, absent additional rules, irreducible subnets would be produced which don’t have a $\lambda$-calculus analogue. As an extreme example, applying an abstraction which doesn’t use its bound variable to an argument which only uses globally-free variables produces a subnet which is disjointed from the root. In order to eliminate all such subnets a final canonicalization reduction step is introduced in $\Delta$-Nets systems with erasure: all parent-child wires starting from the root are traversed and nodes are marked. All non-marked nodes are then erased, and wires that were connected to these nodes are instead connected to erasers. This final canonicalization erasure step dispenses with the need to apply erasure and eraser annihilation rules. In fact, this step can be applied at any point during reduction in order to reduce the net size, effectively trading computation (time) for memory (space). In order to keep memory usage to a minimum, this step should be applied after every application of an abstraction which doesn’t use its bound variable.
In order to ensure that no reduction operations are applied in a subnet that is later going to be erased, a sequential leftmost-outermost reduction order needs to be followed. Therefore, in the $\Delta A$-Nets system, fan annihilations are applied in leftmost-outermost order, with the final erasure canonicalization step ensuring perfect confluence, and producing a normal canonical $\Delta A$-net.
@@
During reduction of $\Delta I$- and $\Delta K$-Nets, replicators can combine into replicator trees. A replicator tree is a subnet containing only replicator agents such that each one’s principal port is connected to an auxiliary port of another, except for the tree’s root replicator. In general, replicators that belong to the same tree cannot be merged during reduction because they may be paired with other replicators elsewhere. However, if consecutive replicators in a replicator tree are known to be unpaired, they can be safely merged together. Unpaired replicator merging is a canonicalization rule present in both $\Delta I$- and $\Delta K$-Nets. Additionally, in the $\Delta K$-Nets system, it is possible to eliminate unpaired replicators’ auxiliary ports which are connected to erasers—a canonicalization rule called unpaired replicator decay. As part of unpaired replicator decay, if an unpaired replicator is left with a single auxiliary port with a level delta of zero, it is replaced by a wire, as it is equivalent to one. This canonicalization rule is applied to all unpaired replicators as part of the erasure canonicalization step. It can also be lazily applied to the involved unpaired replicator, immediately before the fan replication, replicator replication, and aux fan replication rules.
@@
Replicators start as unpaired, and this status is propagated across interactions. When an unpaired replicator interacts with a fan, the status of both resulting replicators changes to unknown. If an unpaired replicator (A) is connected to a consecutive replicator (B) of unknown status via an auxiliary port, and a certain local constraint is met, then the consecutive replicator can be determined to be unpaired, and the two can then be merged. The constraint is met when the second replicator’s level is greater than or equal to the first replicator’s level, but no greater than the first replicator’s level plus the level delta of the auxiliary port that connects them: $0\leq l_{B}-l_{A}\leq d$. Under this constraint, no replicator is able to interact with the second replicator before the first replicator is annihilated. Since the first replicator is unpaired, it can never be annihilated, and the second one must be unpaired as well.
@@
It is possible to limit replicators to have at most two auxiliary ports, or even exactly two auxiliary ports. This imposes the smallest possible upper bound on interaction rule complexity and agent size, making the total number of interactions an effective measure of time complexity, and the total number of agents an effective measure of space complexity. A tree of such replicators can stand in for any replicator with any number of auxiliary ports.
Replicator merging is a canonicalization rule and not an interaction rule because it involves two agents that are connected via ports that aren’t both principal. Merging replicators as early as possible reduces the total number of reductions and the total number of agents, improving space and time efficiency. The reduction order which guarantees that replicator merges happen as early as possible, minimizing the total number of reductions, is a sequential leftmost-outermost order—the optimal reduction order for $\Delta A$-, $\Delta I$- and $\Delta K$-Nets. In fact, any reducible pair that eventually reaches the leftmost-outermost position unchanged can be reduced at any time. For example, all potential annihilations in the spine that don’t involve unpaired replicators can be applied in any order, because they will eventually reach the top of the spine unchanged. On the other hand, a potential commutation that involves an unpaired replicator may not reach the top of the spine unchanged, as the involved replicator may be merged with another beforehand, rendering its early application suboptimal.
To see how leftmost-outermost reduction is optimal in $\Delta I$- and $\Delta K$-Nets, observe that no commutation involving an unpaired replicator can be applied before that replicator is merged—if merging it is at all possible. Take a leftmost-outermost interaction between two replicators. The fan-out replicator is necessarily paired. If the fan-in replicator is unpaired and can eventually be merged, then the replicator merging—or any intermediate reductions leading to it—would occur higher in the spine and would be applied first. This reasoning holds even in the presence of loops. In $\Delta K$-Nets, the leftmost-outermost order is critical not only to achieve optimality but also to ensure that all nets associated with normalizing $\lambda$-terms normalize. As for the remaining rules, the $\Delta$-Nets systems inherit their optimality guarantee from the interaction nets paradigm.
@@
Additionally, in $\Delta I$- and $\Delta K$-Nets, the reduction process needs to be split in two phases. In the first phase the core interaction rules and unpaired replicator merging are applied in leftmost-outermost order until no further reduction can be applied. The reduction process then switches to the second phase, in which the aux fan replication rule (illustrated in Figure 6) replaces the core fan replication rule. This second phase can be alternatively understood as modifying all fans such that the first auxiliary port becomes the principal port. This process transforms the sharing structures so that all fan-out replicators are eliminated, all appropriate subnets are replicated, and all fan-in replicators accumulate at the variable port of abstraction fans. The result is a canonical $\Delta I$-net or, after the final erasure canonicalization step, a canonical $\Delta K$-net. Since all normal $\Delta$-nets are canonical, the $\Delta$-Nets systems are all Church–Rosser confluent.
@@ 5.  Conclusion
The existence of nonlinear proper $\Delta$-nets which are not canonical reflects the additional degrees of freedom that interior sharing introduces, which are not present in the $\lambda$-calculus. A given $\lambda I$-term can potentially be represented by many different proper $\Delta I$-nets, which differ only with respect to their sharing structure. As an example, take a proper $\lambda I$-net $n$, with $\phi_{I}^{-1}(\Omega_{I}(n))=MM$, for some $\lambda I$-term $M$. The $\lambda I$-net $n$ may have two distinct but equal subnets that each represent $M$, or it may have a single such subnet which is shared among the two occurrances in the application. In fact, the sharing structure of $n$ could be arbitrarily complex. These additional degrees of freedom allow $\Delta$-Nets to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is applied which is rendered unneccessary later, and no reduction operation which is necessary is applied more than once.